# Out-of-Distribution Runtime Adaptation with Conformalized Neural Network Ensembles

### v1

@@ V-A Conformal Prediction
To define a valid prediction region, we use the spectral norm of the empirical covariance matrix, given in (7), as the nonconformity score: larger values of $\rho$ signify greater degrees of nonconformity. The following remark results from the application of Lemma $1$ in [40], provided that $\rho^{(i)}$ for $i=1,\dots,N$, the noncomformity measures of the data points in $D_{\mathrm{cal}}$, are exchangeable, with $\rho_{t+1}$ being the nonconformity score associated with the prediction of the pedestrian position at time step $t+1$.
Remark 1.
@@
Given a failure probability ${\delta\in(0,1)}$, a calibration dataset ${D_{\mathrm{cal}}\subset D}$, and the predictions $f_{a}^{i}(\cdot)$${\forall i}$, we have:
$P(\rho_{t+1}\leq C)\geq 1-\delta,$ (9)
for ${t>0}.$
@@
Remark 1 states a standard result from Conformal Prediction that may seem impossible on first reading. Even with very little data, we obtain an exact probability guarantee. For example, with $10$ calibration data points, we can obtain a test with an exact 10% error rate. To resolve this confusion, we must understand that this result considers both the calibration data and the new unseen data points as random variables. In other words, this error rate is obtained exactly when marginalizing over all calibration sets drawn i.i.d. from the data distribution. However, in an engineering application, the calibration data are sampled once and fixed. This results is a probability distribution over the obtained error rate for the OOD classifier, which requires a more sophisticated statistical analysis [32, 2] beyond the scope of this paper. We state the distribution of coverage conditioned on the calibration set in the following theorem.
Theorem 1.
@@
where ${\delta\in(0,1)}$ denotes a specified failure probability.
Proof.
The proof is presented in [44]. Hence, we omit the proof here. ∎
@@
Theorem 1 indicates that the error rate follows a Beta distribution, which can be used to obtain a confidence that the test holds with a given error rate. E.g., we may calibrate for a $10\%$ error rate, but find from this Beta distribution that we will obtain an error rate less than or equal to $10\%$ with $60\%$ probability and at most a $12\%$ error rate with $90\%$ probability. We note that with more samples in our calibration set, we obtain a tighter Beta distribution, resulting in a greater likelihood that we obtain an error rate close to the calibrated value. This more nuanced analysis is discussed in our simulation results in Sec. VII-F, where we calibrate for a $4\%$ error rate, and obtain a $4.4\%$ error rate for our specific calibration set.
@@
We note that naive generation of the calibration dataset would result in the violation of the exchangeability assumption, since trajectories of the dynamic agents collected within the same interaction between the agents and the autonomous robot are not independent. In particular, such trajectories are time-correlated, making them non-exchangeable. To address this issue, we generate the calibration dataset from $N$ trajectories that are independent, obtained from $D$. We partition the $k$ trajectories in $D$ into the training dataset $D_{\mathrm{train}}$, with $N$ trajectories, and the test dataset $D_{\mathrm{test}}$ (with the remaining trajectories), then we form the calibration dataset $D_{\mathrm{cal}}$ by uniformly randomly removing one data point from each trajectory in $D_{\mathrm{train}}$ and placing the relevant data in the calibration set. We note that the resulting calibration dataset satisfies the exchangeability assumption, by satisfying the stronger assumption that trajectories are independent and identically-distributed. To complete the calibration procedure in the conformal prediction framework, we compute the noncomformity measure associated with each data point in $D_{\mathrm{cal}}$, which is used in determining the quantile of the empirical distribution corresponding to the desired value of $\delta$, itself used in the specification of $C$.
VI OOD Adaptive Controller Design
@@
Given an out-of-distribution detector, we compose a safe OOD-adaptive control architecture that utilizes the MPC-based controller in (4) when the dynamic agents in the environment of the autonomous robots are acting nominally, and switches over to a conservative reachable-set-based approach, when out-of-distribution behavior is detected. Algorithm 1 summarizes our proposed Out-of-distribution Adaptive Conformalized Model Predictive Control (OOD-MPC), which is illustrated in Figure 3.
@@
Algorithm 1 OOD-MPC: Out-of-distribution Adaptive Conformalized Model Predictive Control
Figure 3: Architecture of control system: neural network predictions are used for OOD detection, and determination based on OOD detector selects control strategy to be used until next evaluation.
@@
We calibrate the OOD detector prior to solving the MPC problems for the control inputs, given a calibration dataset $D_{\mathrm{cal}}$ and a desired failure probability $\delta$ or equivalently the corresponding index of the sample in the sorted set of samples of nonconformity scores. Here, we describe the calibration process given the value of this index. Given $N$ samples in the sorted calibration dataset $D_{\mathrm{cal}}$ and a desired index value $K$, the probability that a data point randomly sampled from a trajectory in the empirical distribution has a nonconformity score less than that of the $K$th point in $D_{\mathrm{cal}}$ is given by:
@@
provided the samples are exchangeable and assuming randomness over $\rho^{(i)}$${\forall i}$, (see Theorem D.1 in [2]). This corresponds to our desired failure probability $\delta$, as
@@
VII Simulations
In this section, we evaluate the performance of our controller in both in-distribution and out-of-distribution settings, and compare its performance to that of controllers presented in [27] and [22], which utilize conformal prediction to provide probabilistic safety guarantees. We consider the problem of designing a controller that enables an autonomous vehicle to safely navigate to its destination without colliding with a pedestrian crossing the road at a crosswalk. In the in-distribution case, which we refer to as nominal behavior, the pedestrian follows a trajectory from the test dataset, $D_{\mathrm{test}}$. Meanwhile, in the the out-of-distribution case referred to as insurance fraud behavior, the pedestrian actively attempts to force a collision with the autonomous vehicle, by approaching the vehicle while walking at maximum speed. We note that this out-of-distribution behavior constitutes a worst-case scenario where the pedestrian seeks to force the autonomous vehicle to fail to either reach its destination or to avoid collisions. To overcome such an attack, an autonomous vehicle must detect the malicious (out-of-distribution) behavior of the pedestrian and adapt its controller to achieve its goals. In our simulations, the pedestrian in the “insurance fraud” version of each simulation acts nominally for an initial $1.3$ seconds, and thereafter turns to insurance fraud behavior.
VII-A Dynamics Model
The dynamics model of the autonomous vehicle is similar to the Reeds-Shepp car [31], a car that can move both forwards and backwards as well as turn left and right. We denote the state of the autonomous vehicle by ${\mathbf{x}_{t}=(x_{t}^{\intercal},\theta_{t},V_{t},\kappa_{t})^{\intercal}}$, consisting of its $2$-D position ${x_{t}\in\mathbb{R}^{2}}$, its attitude (heading-angle) ${\theta_{t}\in\mathbb{R}}$, its longitudinal velocity ${V_{t}\in\mathbb{R}}$, and the curvature of its trajectory ${\kappa_{t}\in\mathbb{R}}$. We impose limits on the vehicle’s speed with ${\lvert V_{t}\rvert\leq V_{\max}}$${\forall t}$, where $V_{\max}$ denotes the maximum attainable speed by the vehicle in any direction. Further, we constrain the curvature of the vehicle’s trajectory with ${\lvert\kappa_{t}\rvert\leq\kappa_{\max}}$${\forall t}$, where $\kappa_{\max}$ denotes the maximum allowable curvature. The vehicle’s control input at time $t$, denoted by ${\mathbf{u}_{t}\in\mathbb{R}^{2}}$, comprises of its acceleration ${a_{t}\in\mathbb{R}}$ and the pinch (the first derivative of the curvature) of its trajectory ${p_{t}\in\mathbb{R}}$. Given the state and control inputs of the autonomous vehicle, we describe its dynamics in discrete time by the model:

### v3

@@ 5 Out-of-Distribution Detection
To define a valid prediction region, we use the spectral norm of the empirical covariance matrix as the nonconformity score: larger values of $\rho$ signify greater degrees of nonconformity. The following remark results from Lemma $1$ in (Tibshirani et al., 2019), provided that $\rho^{(i)}$ for $i=1,\dots,N$, the noncomformity measures of the data points in $D_{\mathrm{cal}}$, are exchangeable, with $\rho_{t+1}$ being the nonconformity score associated with the prediction of the pedestrian position at time step $t+1$.
Remark 5.1.
@@
Given a failure probability ${\delta\in(0,1)}$, a calibration dataset ${D_{\mathrm{cal}}\subset D}$, and the predictions $f_{a}^{i}(\cdot)$${\forall i}$, we have: ${P(\rho_{t+1}\leq C)\geq 1-\delta},$ for ${t>0},$ where $C$ represents the prediction region associated with $\delta.$
@@
Remark 5.1 states a standard result from Conformal Prediction, which allows strong guarantees despite its simplicity. For example, with $10$ calibration data points, we can obtain a test with an exact 10% error rate. To understand this more easily, note that this result considers both the calibration data and the new unseen data points as random variables. In other words, this error rate is obtained exactly when marginalizing over all calibration sets drawn i.i.d. from the data distribution. However, in an engineering application, the calibration dataset is sampled once and fixed, which results in a probability distribution over the obtained error rate for the OOD classifier, which requires a more sophisticated statistical analysis (Shafer and Vovk, 2008; Angelopoulos and Bates, 2021) beyond the scope of this paper. We state the distribution of coverage conditioned on the calibration set in the following theorem.
Theorem 5.2.
@@
where $N$ is the size of the $D_{cal}$ dataset and $K$ is the index value of the nonconformity score used to set $C$ when the scores in $D_{cal}$ are placed in nondecreasing order.
Proof 5.3.
The proof is presented in (Vovk, 2012). Hence, we omit the proof here.
@@
Theorem 5.2 indicates that the error rate follows a Beta distribution, which can be used to obtain a confidence that the test holds with a given error rate. E.g., we may calibrate for a $10\%$ error rate, but find from this Beta distribution that we will obtain an error rate less than or equal to $10\%$ with $60\%$ probability and at most a $12\%$ error rate with $90\%$ probability. We note that with more samples in our calibration set, we obtain a tighter Beta distribution, resulting in a greater likelihood that we obtain an error rate close to the calibrated value. This more nuanced analysis is discussed in Appendix C.10, where we calibrate for a $4\%$ error rate, and obtain a $4.4\%$ error rate for our specific calibration set. At runtime, we detect OOD behavior when the non-conformity score $\rho_{t+1}$ exceeds the calibrated threshold $C$.
@@
We note that naive generation of the calibration dataset would result in the violation of the exchangeability assumption, since trajectories of the dynamic agents collected within the same interaction between the agents and the autonomous robot are not independent. In particular, such trajectories are time-correlated, making them non-exchangeable. To address this issue, we form the calibration dataset $D_{\mathrm{cal}}$ by uniformly randomly removing one data point from each trajectory in $D_{\mathrm{train}}$ and placing the sampled data in the calibration set. We use the remaining datapoints in the trajectory for training the neural networks. We provide a more detailed discussion of the data-generation procedure in Appendix C.2. We note that the resulting calibration dataset satisfies the exchangeability assumption, by satisfying the stronger assumption that trajectories are independent and identically distributed. To complete the calibration procedure in the conformal prediction framework, we compute the noncomformity measure associated with each data point in $D_{\mathrm{cal}}$ and place them in nondecreasing order, to be used in the specification of $C$ based on the quantile of the empirical distribution corresponding to the desired value of $\delta$.
6 Safe, Out-of-Distribution-Adaptive MPC
@@
Given an OOD detector, we compose a safe OOD-adaptive control architecture that uses the MPC-based controller in (2) with the nominally-acting, dynamic agents, and switches over to a conservative reachable-set-based approach when OOD behavior is detected. Algorithm 1 summarizes our proposed method SODA-MPC: Safe, Out-of-distribution-Adaptive Model Predictive Control, which is illustrated in Figure 2 and described in greater detail in Appendix B.
7 Simulations
We evaluate the performance of our controller in both in-distribution and OOD settings, and compare its performance to that of controllers presented in (Muthali et al., 2023) and (Lindemann et al., 2023), which utilize conformal prediction to provide probabilistic safety guarantees, in a problem where an autonomous vehicle attempts to safely navigate to its destination without colliding with a pedestrian crossing the road at a crosswalk. In the in-distribution case, which we refer to as nominal behavior, the pedestrian follows a trajectory from the test dataset, $D_{\mathrm{test}}$. Meanwhile, in the OOD case referred to as insurance fraud behavior, the pedestrian attempts to force a collision by approaching the vehicle at maximum speed, which constitutes a worst-case scenario, after initially acting nominally for $1.3$ seconds. To overcome such an attack, an autonomous vehicle must detect the OOD behavior quickly and adapt its controller to avoid a collision.
@@ 7 Simulations
Out-of-distribution-Adaptive
Model Predictive Control
Algorithm 1 SODA-MPC: Safe,
Out-of-distribution-Adaptive
Model Predictive Control
Figure 2: Architecture of SODA-MPC: SODA-MPC detcts OOD behavior from online observations to select a safe control strategy, depending on the behavior of other agents.
7.1 Small, Feedforward MLP Trajectory-Prediction Models
Here, we examine the effectiveness of SODA-MPC in detecting OOD behavior, in problems where the trajectory-prediction model consists of small, feedforward MLPs. Specifically, we predict the future state of a pedestrian using an ensemble of identical MLPs, each with 66 parameters in total, trained on the VCI dataset (Yang et al., 2019). We present the architecture of these models and the associated training dataset in Appendix C.
Empirical Results. We assess the performance of our control algorithm SODA-MPC against four baselines, namely: (a) SPDE (Lindemann et al., 2023); (b) MARC (Muthali et al., 2023); (c) Reachable Sets Only, a variant of SODA-MPC utilizing only the MPC II controller; and (d) Ensembles Only, a variant of SODA-MPC utilizing only the MPC I controller. We examine each algorithm in $20$ simulations, presenting the results achieved by each algorithm in the subsequent discussion. Each simulation terminates in one of three ways: (a) the autonomous vehicle goes past the pedestrian while avoiding collisions, (b) the autonomous vehicle does not go past the pedestrian but avoids a collision (e.g., when stopping to allow the pedestrian to finish crossing the road), or (c) the autonomous vehicle collides with the pedestrian. In SODA-MPC, we utilize a failure probability of approximately $4\%$ in the conformal prediction procedure for computing prediction regions, associated with ${C=0.012}$. In contrast, the uncertainty quantification for the MARC algorithm (Muthali et al., 2023) corresponds to a failure probability of $5\%$ and the ball-shaped confidence areas for the SPDE algorithm (Lindemann et al., 2023) correspond to a failure probability of $5\%$. In all the figures, we depict the past trajectory of the autonomous vehicle in blue when the autonomous vehicle detects nominal behavior and in red when the autonomous vehicle detects OOD behavior. We depict the past trajectory of the pedestrian in cyan. In Figure 4, we show that SODA-MPC and the more conservative methods MARC and Reachable Sets Only do not collide with the pedestrian, even when the pedestrian attempts to force a collision. In the nominal case, the Ensembles Only method achieves the highest success rate at $70\%$, followed jointly by SODA-MPC and SPDE, which achieve a success rate of $60\%$. Consequently, SODA-MPC provides a desirable tradeoff across both behavior modes.
Figure 3: In a pedestrian-crossing scenario, only SODA-MPC (our method) both passed the horizontal position of the pedestrian in the nominal case and avoided collisions with the pedestrian in the insurance fraud case, compared to the other baselines.
Figure 4: In the top-two rows, the Ensembles Only method avoids collisions in the nominal case (top) and collides in the insurance fraud setting (bottom). In the bottom-two rows, the overly-conservative Reachable Sets Only method avoids collisions in both settings.
While the Ensembles Only method could reliably avoid collisions in the nominal case, the method failed to avoid collisions in the insurance fraud case, since the predictions of the pedestrian’s trajectory differed wildly from the actual trajectory taken by the pedestrian, as depicted in Figure 4. In contrast, the Reachable Sets Only method avoids collisions in all settings; however, the autonomous vehicle was consistently unable to reach its desired goal location during the simulation, as illustrated in Figure 4. We note that the MARC algorithm achieved similar conservative results to those of the Reachable Sets Only method. Although SPDE remains safe while navigating the vehicle to its goal in the nominal case, SPDE does not avoid collisions when the pedestrian attempts to forcefully collide with the vehicle: this can be attributed to the erroneous predictions generated by the neural network ensemble. Though the method adds a buffer around the position of the pedestrian based on the estimated uncertainty of its prediction, the uncertainty is calibrated under the assumption that the true pedestrian trajectory is drawn from the same distribution as in the training data, while in the insurance fraud setting, it is drawn from an altogether different distribution. In contrast, SODA-MPC avoids collisions in all settings, e.g., in the insurance fraud setting, the controller switches from the nominal control strategy to the reachable-set-based control strategy after detecting OOD behavior, depicted in Figure 1. We provide empirical results on the statistical guarantees of our OOD detector in Appendix C.10, showing that our detector achieves a false-positive rate of 4.4% while correctly identifying OOD behavior 91.3% of the time.
7.2 Trajectron++, a Large-Scale Trajectory-Prediction Model
Here, we evaluate the performance of our proposed OOD detector applied to the large-scale trajectory predictor Trajectron++ (Salzmann et al., 2020) as the trajectory predictor on the nuScenes (Caesar et al., 2020) dataset. We provide additional setup details in Appendix C.11.
Empirical Results. For the calibration procedure, we set $K=97$ for $N=100$ for the calibration set point, yielding a detection threshold of $C=0.933$. Here, predictions are evaluated at the trajectory level, rather than at individual time steps: a trajectory as a whole is considered to be anomalous if the uncertainty value for at least one of its constituent time steps is greater than the detection threshold, and is considered to be nominal otherwise. Nominal cases cover 6230 trajectories of pedestrians in the dataset, and OOD cases use 230 trajectories with shifted pedestrian behavior. We use a calibrated error rate of $\delta\approx 4\%$ for individual data points. In this setting, our OOD detector correctly identifies 5361 (86.1%) of the nominal trajectories as nominal, with a false-positive rate of 13.9% (869 trajectories). Likewise, our OOD detector correctly identifies 211 (91.7%) of the insurance fraud trajectories as OOD, with a false-negative rate of 8.3% (19 trajectories), summarized in Table 2 in Appendix C.11. Compared to an ensemble of identically trained MLP models, the increase in uncertainty seen between different modes when an agent was displaying anomalous behavior was still detectable but less pronounced: we believe this may be ascribed to the cost function of the Trajectron++ model being specifically designed to promote divergence between the prediction modes, rather than training them all to react identically to training data points. This phenomenon was counteracted by evaluating predictions at the trajectory level, which increases the overall sensitivity of the system to detections of anomalous data points. As trajectories displaying insurance fraud behavior still have a large number of data points below the detection threshold, this vastly improved the detection rate for anomalous behavior, although the increased sensitivity did increase the incidence of nominal data points being mischaracterized as anomalous compared to the experiments with ensembled feedforward MLP models.
8 Conclusion, Limitations, and Future Work
We introduce SODA-MPC, a safe, adaptive controller that enables autonomous robots to safely navigate in their environments in nominal or OOD settings. SODA-MPC is comprised of a nominal controller based on learned trajectory predictors for in-distribution settings, and safe-by-design controller based on reachable sets for OOD cases. Our OOD detector uses conformal prediction to provide provable probabilistic guarantees on the validity of the prediction regions. We demonstrate the safety and efficacy of our controller in a pedestrian-crossing scenario. With SODA-MPC, the autonomous vehicle avoids collisions with the pedestrian in all settings, and is able to pass the position of the pedestrian when the pedestrian acts in-distribution; whereas some other existing methods either result in collisions, when the pedestrian attempts to force them, or are overly conservative (failing to reach the goal). Our OOD-controller may be too conservative in some settings, especially in dense (congested) scenarios. In future work, we seek to explore the utilization of deep-learned controllers in combination with (or in lieu of) MPC-based controllers in our control framework, which could enable the development of more expressible control frameworks. Further, we are interested in exploring the use of our method in uncertainty quantification to identify and search for areas of the predictor input space not appropriately covered by training data, and in examining other uncertainty quantification metrics for OOD detection.
[…]
This work was supported in part by ONR grant N00014-23-1-2354, NASA ULI grant 80NSSC20M0163, DARPA grant HR001120C0107, and NSF grant FRR 2342246. Toyota Research Institute provided funds to support this work.
Appendix A Conformal Prediction
Conformal prediction was introduced in (Vovk et al., 2005) and (Shafer and Vovk, 2008) to compute prediction regions for complex predictive models without making assumptions about the distribution underlying the predictive model (Angelopoulos and Bates, 2021). We present a brief introduction based on the assumption that the observations arise from independent and identical distributions for easier presentation. However, we note that the theory of conformal prediction only requires the assumption of exchangeability. Observations in a set ${\{\rho^{(1)},\ldots,\rho^{(N)}\}}$ are exchangeable if all possible orderings of these observations are equally likely to occur. Consequently, exchangeability represents a weaker assumption compared to the assumption of independent, identically-distributed observations.
[…]
Appendix B OOD-Adaptive Controller Design
@@
We calibrate the OOD detector prior to solving the MPC problems for the control inputs, given a calibration dataset $D_{\mathrm{cal}}$ and a desired failure probability $\delta$, or equivalently, the corresponding index of the sample in the sorted set of samples of nonconformity scores. Here, we describe the calibration process given the value of this index. Given $N$ samples in the sorted calibration dataset $D_{\mathrm{cal}}$ and a desired index value $K$, the probability that a data point randomly sampled from a trajectory in the empirical distribution has a nonconformity score less than that of the $K$th point in $D_{\mathrm{cal}}$ is given by:
@@
provided the samples are exchangeable and assuming randomness over $\rho^{(i)}$${\forall i}$, (see Theorem D.1 in (Angelopoulos and Bates, 2021)). This corresponds to our desired failure probability $\delta$, as
@@
Appendix C Simulation Setup
C.1 Predictor Architecture
C.2 Dataset
We use the Vehicle-Crowd Intraction[sic] (VCI) dataset (Yang et al., 2019) to train and evaluate the performance of each method in simulation. The dataset used, a subset of the total VCI dataset, consists of $110$ trajectories of pedestrians crossing a roadway, with each trajectory comprised of the position data of a pedestrian for $154$ time steps. In our evaluations, we randomly initialize the position of the pedestrian, relative to the ego vehicle. The horizontal starting positions are sampled from a normal distribution with a mean $40$ meters from the horizontal starting position of the ego vehicle and a standard deviation of $2.5$ meters. The starting vertical position of the pedestrian is determined by the direction of travel in the specific trajectory obtained from the dataset. To generate the training data, we process the dataset into input-output data pairs, where the input consists of the positions of the pedestrian over $14$ consecutive time steps and the output consists of the position of the pedestrian at the following time step.
Of the 110 trajectories, 10 are selected randomly to serve as the test dataset $D_{\mathrm{test}}$ and used in evaluating the performance of each algorithm. For the remaining 100 trajectories, one data pair from each trajectory is selected uniformly randomly and removed, and these 100 data points form the calibration dataset $D_{\mathrm{cal}}$ used in conformal prediction. The remaining data points comprise the training dataset $D_{\mathrm{train}}$, used in training the parameters of each neural network in the ensemble through Adam optimization. These neural networks were used in testing for nearly all of the methods and benchmarks, with the exception of the testing for benchmarks based on (Lindemann et al., 2023) and (Muthali et al., 2023). For the benchmark based on (Lindemann et al., 2023), 20 uniformly randomly selected trajectories from the 100 full trajectories were used to form the calibration dataset, with the remaining 80 trajectories used in training the neural networks. For the benchmark based on (Muthali et al., 2023), no calibration dataset was formed, and all input-output data points from the 100 trajectories were used for training.
One of the limitations of the dataset is the amount of data available: with only 15400 data points in total available across training, calibration and test datasets, it is not possible to make statistical determinations that depend on large amounts of data. To remedy this, a neural network more sophisticated than those used for the predictor ensemble was trained on the existing trajectories and used to generate additional pedestrian trajectories, which we used for the calibration process and statistical analysis. The architecture of the network used to generate new data can be found in Figure 9, where the neural network uses an LSTM layer with 128 neurons, a dense layer of 128 neurons and ReLU activations to transform rectilinear trajectories of 2-dimensional states into trajectories of two-dimensional states mimicking the behavior of a pedestrian.
C.3 OOD Detector
In our simulations, we utilize ${K=97}$, for $N=100$ (i.e., $\delta\approx 3.9604\%$) in calibrating the OOD detector, which yielded good results for reliable OOD detection, with $C$ set to the value of $\rho^{(K)}$ rounded up to the third decimal point. Careful readers will notice that the probabilistic guarantees are specifically placed on the probability of correctly determining that a given data point was generated by the same distribution that generated the training data, but as we will show in the section detailing the results of our experiments, over the course of our experiments we were able to reliably detect changes in the underlying data-generating distribution.
In our simulations, our OOD detector classifies any data point with an uncertainty score above $C$ as being out-of-distribution. We note, however, that the simplicity of this detection framework can result in incorrect classifications of in-distribution samples as OOD, given the probabilistic guarantees in Theorem 5.2. Nevertheless, we can address this challenge using existing methods such as (Bates et al., 2023), which enables the detector to ignore occasional high-uncertainty scores from in-distribution points, but still respond appropriately to persistent high-uncertainty scores that indicate a distribution shift. Further, other uncertainty metrics in anomaly detection, such as those in (Sharma et al., 2021) and (Laxhammar and Falkman, 2011) can be applied in OOD detection. However, these methods require access to the training and calibration data, unlike our method.
C.4 Dynamics Model
[…]